what volume (in L) of 4.35 M CHI solutions is needed to react with 18.5 g of Ca(OH)2


1
Expert's answer
2022-04-29T12:30:10-0400

For a reaction to be completed, the two reagents should have the same amount of moles:

"n(HI) = n(Ca(OH)_2)"

"n(Ca(OH)_2) = \frac{m}{M} = \frac{18.5g}{74.1 g/mol} = 0.25 mol"

The mole number of HI is 0.25 mol according to the first equation and can be written in the following way:

"n(HI) = c\times V"

Now the volume is:

"V(HI)=\frac{n}{c}=\frac{0.25 mol}{4.35 mol/l} = 0.057 L"